# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
Since $u_{1},w_{3}\notin V_{\rm bad}$, there exist $i_{0},i_{1}\in\mathcal{C}_{1}\setminus{\rm col}(P^{3})$ such that $u_{1}\in A_{i_{0}}$ and $w_{3}\in B_{i_{1}}$. By the construction of $P^{3}$, one may assume that $v\not\in V(P^{3})$ unless $v\in V_{\rm bad}\cup X^{\prime}$, and $c\not\in{\rm col}(P^{3})$ unless $c\in\mathcal{C}_{\rm bad}$. Similar to the above two cases, we know that each digraph $D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{2}$ can be viewed as an undirected graph with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$; each digraph $D_{i}[A\setminus V(P^{3})]\cup D_{i}[B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{1}$ can be viewed as the union of two disjoint undirected graphs with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$. In what follows, we proceed by considering the parity of $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$.
Subcase 3.1. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is odd.
Split $A\setminus V(P^{3})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$, and split $B\setminus V(P^{3})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|+1}{2}$. Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus{\rm col}(P^{3})\}$. By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $x\in I_{1}$ and $y\in J_{1}$. Since $x,y\not\in V_{\rm bad}$, there exist $j_{0},j_{1}\in\mathcal{C}_{1}$ such that $x\in A_{j_{0}}$ and $y\in B_{j_{1}}$.
Suppose that $|I_{2}|=|A\setminus V(P^{3})|-|I_{1}|=k_{1}$ and $|J_{2}|=|B\setminus V(P^{3})|-|J_{1}|=k_{2}$. We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$. Then split the color set $\mathcal{C}_{1}\setminus({\rm col}(P^{3})\cup\{j_{0},j_{1}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$. Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$. Applying Claim 5.5 again yields that there is a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{i_{0}}}^{-}(u_{1},I_{2}^{\prime}),\,y_{a}\in N_{{j_{0}}}^{+}(x,I_{2})$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{i_{1}}}^{+}(w_{3},J_{2}^{\prime}),\,y_{b}\in N_{{j_{1}}}^{-}(y,J_{2})$. Thus, $u_{1}P^{3}w_{3}x_{b}Q_{b}y_{b}yQxy_{a}Q_{a}x_{a}u_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$, a contradiction.
Subcase 3.2. $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even.
In this subcase, we first prove the following claim.
@@
Split $A\setminus V(Q_{1}\cup Q_{2})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$, and split $B\setminus V(Q_{1}\cup Q_{2})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$. Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{4}\})\}$. By Claim 5.5, there exists a transversal undirected path $Q$ inside $\mathcal{D}_{1}$ with endpoints $z_{1}^{\prime\prime}\in N_{{\ell_{4}}}^{+}(z_{1}^{\prime})\cap J_{1}$ and $y\in I_{1}$. Since $y\not\in V_{\rm bad}$, there exists $\ell_{5}\in\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3}\})$ such that $y\in A_{\ell_{5}}$.
@@
Assume that $|I_{2}|=k_{1}$ and $|J_{2}|=k_{2}$. We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, and split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$. Split the color set $\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3},\ell_{5}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$. Let $\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}$ and $\mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}$. By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with endpoints $x_{a}\in N_{{\ell_{1}}}^{-}(z_{1})\cap I_{2}^{\prime},\,y_{a}\in N_{{\ell_{5}}}^{+}(y)\cap I_{2}$; and a transversal undirected path $Q_{b}$ inside $\mathcal{D}_{b}$ with endpoints $x_{b}\in N_{{\ell_{2}}}^{+}(z_{k})\cap J_{2}^{\prime},\,y_{b}\in N_{{\ell_{3}}}^{-}(z_{t}^{\prime})\cap J_{2}$. Hence, $z_{1}Q_{1}z_{k}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}z_{1}^{\prime\prime}Qyy_{a}Q_{a}x_{a}z_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$. ∎
@@
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $|A|\leq|B|$, each vertex in $A$ has an in-neighbor in $B$ in each $D_{i}$ with $i\in\mathcal{C}_{1}\cup{\rm col}(M)$. Suppose that there exists an edge $y_{2}y_{1}\in E(D_{i_{2}}[B,A])$ for some $i_{2}\in\mathcal{C}_{1}\cup{\rm col}(M)$, where $y_{1}\in A\setminus(V(P^{3})\cup V_{\rm bad})$ and $y_{2}\in B$. If $i_{2}\in{\rm col}(M)$, then assume ${uv}\in E(M[A])\cap E(D_{i_{2}})$. By Claim 5.3 (i)-(ii), we connect $P^{2}$, all rainbow edges in $M-uv$ and all rainbow edges in $M^{\prime}$ in turn. This results a rainbow directed path $\tilde{P}^{3}$ with endpoints $\tilde{u}_{1}\in A\setminus V_{\rm bad}$ and $\tilde{w}_{3}\in B\setminus V_{\rm bad}$. Therefore, $P^{3}$ can be written as $uvv^{\prime}\tilde{u}_{1}\tilde{P}^{3}\tilde{w}_{3}$, where the colors of ${vv^{\prime}}$ and ${v^{\prime}\tilde{u}_{1}}$ are in $\mathcal{C}_{2}$. Since $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even, one has $|\mathcal{C}_{2}\setminus{\rm col}(\tilde{P}^{3})|$ is even. Hence one may let $P^{3}:=\tilde{P}^{3}$ if $i_{2}\in{\rm col}(M)$ and further assume $i_{2}\not\in{\rm col}(P^{3})$.
@@
If $y\in X^{\prime}$, then by adjusting the selections of $y_{2}^{1},y_{2}^{2}$ and $j_{1},j_{2}$, one may get a rainbow directed path $P^{3}$ such that $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd. This can be reduced to Subcase 3.1.
If $y\in V_{\rm bad}$, then $|\mathcal{C}_{2}\setminus{\rm col}(\hat{P^{3}})|$ is even. Let $\hat{P}:=y_{2}^{1}y_{2}y_{1}$ if $y_{2}^{1}\in B$, and let $\hat{P}:=y^{\prime}y_{2}^{1}y_{2}y_{1}$ with $y^{\prime}y_{2}^{1}\in E(D_{c^{\prime}})$ for unused color $c^{\prime}\in\mathcal{C}_{2}$ and unused vertex $y^{\prime}\in B$ if $y_{2}^{1}\in A$. Consequently, $\hat{P^{3}}$ and $\hat{P}$ are two rainbow directed paths satisfying all conditions in Claim 5.12. Hence, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. ∎
In order to complete the proof of Theorem 5.1, we need to show the existence of transversal directed Hamilton cycles in Step 1 of Case 1. We first state a preliminary result from [11].
Lemma 5.13 ([11]).
Let $\mathcal{C}$ be a set of colors, and $\mathcal{G}=\{G_{i}[Y,B]:i\in\mathcal{C}\}$ be a collection of bipartite graphs with the common bipartition $Y\cup B$ such that $7|Y|<|B|\leq\frac{3}{5}|\mathcal{C}|$. If $\sum_{i\in\mathcal{C}}|E(D_{i}[Y,B])|\geq t|B||\mathcal{C}|$ for some integer $t$ with $1\leq t\leq|Y|$, then $\mathcal{G}$ contains $t$ disjoint rainbow star on $5$ vertices, each with the center in $Y$ and other vertices in $B$.
Lemma 5.14.
Proof.
$=(|A|-\delta^{\frac{1}{2}}|Y_{1}|)|B||\mathcal{C}^{\prime}|.$ (4)
It implies that $|Y_{1}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. Similarly, $|Y_{2}|\leq\frac{\gamma n}{\sqrt{\delta}}\leq 3\sqrt{\delta}n$. For a vertex $v\in Y_{1}\setminus Y_{2}$ (resp. $v\in Y_{2}\setminus Y_{1}$), we have $d_{i}^{-}(v,B)>(1-\delta^{\frac{1}{4}})|B|$ (resp. $d_{i}^{+}(v,B)>(1-\delta^{\frac{1}{4}})|B|$) for at least $(1-\delta^{\frac{1}{4}})|\mathcal{C}^{\prime}|$ colors $i\in\mathcal{C}^{\prime}$. Notice that Claim 5.5 and Claim 5.3 (i)-(ii) hold by setting $V_{\rm bad}:=Y_{1}\cup Y_{2}$ and $\mathcal{C}_{2}=\mathcal{C}^{\prime}$.
Case 1. $|Y_{1}|,|Y_{2}|\geq\gamma n$.
In view of Lemma 5.13, there exists a subset $Y_{1}^{\prime}\subseteq Y_{1}$ with size $\gamma n$ such that vertices in $Y_{1}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[Y_{1},B]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{1}\setminus Y_{1}^{\prime}$; and a subset $Y_{2}^{\prime}\subseteq Y_{2}$ with size $\gamma n$ such that vertices in $Y_{2}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow $P_{3}$ inside $\{D_{i}[B,Y_{2}]:i\in\mathcal{C}^{\prime}\}$ with centers in $Y_{2}\setminus Y_{2}^{\prime}$. Recall that each vertex in $Y_{1}$ (resp. $Y_{2}$) has at least $\left\lceil\frac{n}{2}\right\rceil-(1-\delta^{\frac{1}{4}})|B|>4|Y|$ out-neighbors (resp. in-neighbors) in $A$ for at least $\delta^{\frac{1}{4}}|\mathcal{C}^{\prime}|>\delta^{\frac{1}{4}}(1-\delta)n>4|Y|$ digraphs $D_{i}$ with $i\in\mathcal{C}^{\prime}$. Therefore, by using colors in $\mathcal{C}^{\prime}$,
  • vertices in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $(Y_{1}\cup Y_{2})\setminus(Y_{1}^{\prime}\cup Y_{2}^{\prime})$ and endpoints in $B$,
  • vertices in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $B$ and ending at $A$ with centers in $Y_{1}^{\prime}\setminus Y_{2}^{\prime}$,
  • vertices in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ starting at $A$ and ending at $B$ with centers in $Y_{2}^{\prime}\setminus Y_{1}^{\prime}$,
  • vertices in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ can be covered by disjoint rainbow directed $P_{3}$ with centers in $Y_{1}^{\prime}\cap Y_{2}^{\prime}$ and endpoints in $A$.
For each $y\in Y_{1}\cup Y_{2}$, denote the rainbow directed $P_{3}$ with center $y$ by $P_{y}:=y^{1}yy^{2}$ with colors $c_{y}^{1}$ and $c_{y}^{2}$. Let $\mathbf{P}=\{P_{y}:y\in Y_{1}\cup Y_{2}\}$. By using Claim 5.3, one may connect all of the above rainbow directed $P_{3}$ into a single rainbow directed path $P^{1}$ starting at $A$ and ending at $B$. Then
@@
Greedily choose two disjoint rainbow directed paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)\}$ with lengths $\lceil\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rceil$ and $\lfloor\frac{|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|}{2}\rfloor$ respectively, such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$. Applying Claim 5.3 (i) to connect $P^{1}$, all rainbow edges in $M$, $P_{A}$ and $P_{B}$ in turn, we get a single rainbow path $P^{2}$ with endpoints in different parts. Clearly, $|A\setminus V(P^{2})|=|B\setminus V(P^{2})|$. Together with Claim 5.5, we know $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired.
Subcase 1.2. $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|$ have the different parity.
Notice that $|B|=\lfloor\frac{n}{2}\rfloor-\gamma n$. Hence for each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $B$ has at least $\gamma n+1$ in-neighbors in $A$. Therefore, each vertex in $B$ has at least one in-neighbor outside $Y_{1}^{\prime}$. We first consider $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|\geq 1$. Choose an unused vertex $y\in B$ and a color $c\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)$. Suppose that ${xy}\in E(D_{c}[A,B])$ with $x\notin Y_{2}^{\prime}$.
$\bullet$ $x\in A\setminus(Y_{1}\cup Y_{2})$. Then $x$ can be chosen outside $V(P^{1}\cup M)$. By Claim 5.3 (i)-(ii), one may connect $xy$ and $P^{1}$ into a single rainbow directed path $P$. Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
$\bullet$ $x\in(Y_{1}\cup Y_{2})\setminus Y_{1}^{\prime}$. Then replace the end vertex of $P_{x}$ by $y$. Applying a similar discussion as Subcase 1.1, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
Therefore, it suffices to consider $n$ is odd and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(M)|=0$. For each $D_{i}$ with $i\in\mathcal{C}$, each vertex in $A$ has at least one in-neighbor outside $Y_{1}^{\prime}$. Choose an unused vertex $y\in A\setminus(Y_{1}\cup Y_{2})$ and a color $c\in\mathcal{C}$. Then there exists a vertex $x\in A\setminus Y_{1}^{\prime}$ such that ${xy}\in E(D_{c}[A])$. By a similar discussion as above, we obtain a transversal directed Hamilton cycle inside $\mathcal{D}$, as desired.
Case 2. At least one of $Y_{1}$ and $Y_{2}$ has size less than $\gamma n$.
In this case, we consider only $|Y_{1}|\geq\gamma n$ and $|Y_{2}|<\gamma n$; the other two cases can be discussed similarly. Based on the proof of Case 1, we know that all vertices in $Y_{1}$ can be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}$ and endpoints in $A\setminus(Y_{1}\cup Y_{2})$. Furthermore, each vertex in $Y_{2}\setminus Y_{1}$ can be covered by a rainbow directed $P_{3}$ with center in $Y_{2}\setminus Y_{1}$ and start vertex in $A\setminus(Y_{1}\cup Y_{2})$ using colors in $\mathcal{C}^{\prime}$. (In this process, for the choice of the other endpoint of each rainbow directed $P_{3}$, we prefer using vertices in $A$ over vertices in $B$.) By using unused colors in $\mathcal{C}$ and unused vertices in $A\setminus Y$, one may extend those rainbow directed $P_{3}$ as long as possible or choose all other disjoint rainbow directed paths. Let $\mathbf{P}=\{Q_{1},Q_{2},\ldots,Q_{t}\}$ be a set consisting of all disjoint rainbow directed paths in the above, where each $Q_{i}$ has at least one endpoint in $A\setminus Y$ and length $s_{i}\ (1\leq i\leq t)$ inside $A$.
Subcase 2.1. $|B|\geq|A|-(s_{1}+\cdots+s_{t})-\sigma$.
In this subcase, there exists a set $\mathbf{P}^{\prime}=\{Q_{1}^{\prime},Q_{2}^{\prime},\ldots,Q_{\ell}^{\prime}\}$ such that $|B|=|A|-(s_{1}^{\prime}+\cdots+s_{\ell}^{\prime})-\sigma$, the endpoints of $Q_{i}^{\prime}$ are not in $Y$ for all $i\in[\ell]$ and $Y_{2}\cup Y_{1}^{\prime}\subseteq V(Q_{1}^{\prime}\cup\ldots\cup Q_{\ell}^{\prime})$, where $s_{i}^{\prime}:=|E(Q_{i}^{\prime}[A])|$ for each $i\in[\ell]$. By Claim 5.3 (i)-(ii), we connect all rainbow directed paths in $\mathbf{P}^{\prime}$ into a single rainbow directed path $P^{1}$, whose endpoints are in different parts. Since $|A|-|B|=2\gamma n+\sigma$, we have $\ell\leq s_{1}^{\prime}+\cdots+s_{\ell}^{\prime}=2\gamma n$. Therefore, $|E(P^{1})|\leq 8\gamma n$ and $|A\setminus V(P^{1})|-\sigma=|B\setminus V(P^{1})|$.
Recall that all vertices in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ can also be covered by a set of disjoint rainbow directed $P_{3}$ with centers in $Y_{1}\setminus V(\mathbf{P}^{\prime})$ and both endpoints in $B$. Applying Lemma 5.3 (i) again to connect those rainbow directed paths with $P^{1}$, we get a rainbow directed path $P^{2}$ with length at most $8\gamma n+12\sqrt{\delta}n$, and with endpoints in different parts.
[…]
Based on Claim 5.3 (i), one may connect $P^{1}$ and all rainbow directed edges in $M$ into a single rainbow directed path $P^{2}$, whose endpoints are in different parts. Clearly, $|E(P^{2})|\leq 8\gamma n+4{\delta}n$. Next, we are to choose two disjoint rainbow paths $P_{A}$ and $P_{B}$ inside $\{D_{i}[A\setminus V(P^{1}\cup M)]\cup D_{i}[B\setminus V(P^{1}\cup M)]:i\in\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)\}$ such that $V(P_{A})\subseteq A$ and $V(P_{B})\subseteq B$ respectively, whose lengths are determined by the parity of $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$. Clearly, if $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have the same parity, then by Claim 5.5, we know $\mathcal{G}$ contains a transversal Hamilton cycle, a contradiction. Hence $n$ and $|\mathcal{C}^{\prime\prime}\setminus{\rm col}(P^{1}\cup M)|$ have different parity. By similar arguments as the proof of Step 4 in Case 1, we can prove our result.
Subcase 2.2. $|B|<|A|-(s_{1}+\cdots+s_{t})-\sigma$.
5.2 Proof of Theorem 5.2
In this subsection, we give the proof of Theorem 5.2.
@@
Recall that if $|\mathcal{C}_{3}|\geq 2\sqrt{\delta}n$, then $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|<2\sqrt{\delta}n$. Denote $[m]:=\mathcal{C}_{3}\setminus I^{2}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$. By adding colors to $\mathcal{C}_{\rm bad}$ if necessary, we may assume $m=(1-4\sqrt{\delta})n$. By (B2), we have $|C_{1}^{k}\Delta C_{i}^{k}|\leq 2\delta n$ for all $k\in[4]$ and all $i\in[m]$. Expand $C_{1}^{1}\cup C_{1}^{2}\cup C_{1}^{3}\cup C_{1}^{4}$ into a partition $C^{1}\cup C^{2}\cup C^{3}\cup C^{4}$ of $V$ such that $|C^{1}|-|C^{3}|\leq 1$ and $|C^{2}|=|C^{4}|$. Hence, for $i\in[2,m]$ and $k\in[4]$, we have $|C^{k}\Delta C_{i}^{k}|\leq 2\delta n+2{\epsilon}n<3\delta n$.

### v2

@@ 1. $(\dagger)$
Since $u_{1},w_{3}\notin V_{\rm bad}$, there exist $i_{0},i_{1}\in\mathcal{C}_{1}\setminus{\rm col}(P^{3})$ such that $u_{1}\in A_{i_{0}}$ and $w_{3}\in B_{i_{1}}$. By the construction of $P^{3}$, one may assume that $v\not\in V(P^{3})$ unless $v\in V_{\rm bad}\cup X^{\prime}$, and $c\not\in{\rm col}(P^{3})$ unless $c\in\mathcal{C}_{\rm bad}$. Similar to the above two cases, we know that each digraph $D_{i}^{\pm}[A\setminus V(P^{3}),B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{2}$ can be viewed as an undirected graph with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$; each digraph $D_{i}[A\setminus V(P^{3})]\cup D_{i}[B\setminus V(P^{3})]$ with $i\in\mathcal{C}_{1}$ can be viewed as the union of two disjoint undirected graphs with minimum degree at least $(\frac{1}{2}-45\sqrt{\delta})n$. We first prove the following claim.
@@
Split $A\setminus V(Q_{1}\cup Q_{2})$ into $I_{1}$ and $I_{2}$ such that $|I_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$, and split $B\setminus V(Q_{1}\cup Q_{2})$ into $J_{1}$ and $J_{2}$ such that $|J_{1}|=\frac{|\mathcal{C}_{2}\setminus{\rm col}(Q_{1}\cup Q_{2})|}{2}$. Let $\mathcal{D}_{1}:=\{D_{i}^{\pm}[I_{1},J_{1}]:i\in\mathcal{C}_{2}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{4}\})\}$. By Lemma 6.2, $\mathcal{D}_{1}$ contains a transversal undirected path $Q$ with endpoints $z_{1}^{\prime\prime}\in N_{{\ell_{4}}}^{+}(z_{1}^{\prime})\cap J_{1}$ and $y\in I_{1}$. Since $y\not\in V_{\rm bad}$, there exists $\ell_{5}\in\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3}\})$ such that $y\in A_{\ell_{5}}$.
@@
Assume that $|I_{2}|=k_{1}$ and $|J_{2}|=k_{2}$. We further split $I_{2}$ into $I_{2}^{\prime}$ and $I_{2}^{\prime\prime}$ such that $|I_{2}^{\prime}|=\lceil\frac{k_{1}}{2}\rceil$, and split $J_{2}$ into $J_{2}^{\prime}$ and $J_{2}^{\prime\prime}$ such that $|J_{2}^{\prime}|=\lceil\frac{k_{2}}{2}\rceil$. Split the color set $\mathcal{C}_{1}\setminus({\rm col}(Q_{1}\cup Q_{2})\cup\{\ell_{1},\ell_{2},\ell_{3},\ell_{5}\})$ into two subsets $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ such that $|\mathcal{C}_{a}|=k_{1}-1$ and $|\mathcal{C}_{b}|=k_{2}-1$. Let
$\mathcal{D}_{a}:=\{D_{i}^{\pm}[I_{2}^{\prime},I_{2}^{\prime\prime}]:i\in\mathcal{C}_{a}\}\ \ \text{and}\ \ \mathcal{D}_{b}:=\{D_{i}^{\pm}[J_{2}^{\prime},J_{2}^{\prime\prime}]:i\in\mathcal{C}_{b}\}.$
Applying Lemma 6.2 again yields that
  • $\mathcal{D}_{a}$ contains a transversal undirected path $Q_{a}$ with endpoints $x_{a}\in N_{{\ell_{1}}}^{-}(z_{1})\cap I_{2}^{\prime},\,y_{a}\in N_{{\ell_{5}}}^{+}(y)\cap I_{2}$;
  • $\mathcal{D}_{b}$ contains a transversal undirected path $Q_{b}$ with endpoints $x_{b}\in N_{{\ell_{2}}}^{+}(z_{k})\cap J_{2}^{\prime},\,y_{b}\in N_{{\ell_{3}}}^{-}(z_{t}^{\prime})\cap J_{2}$.
Hence, $z_{1}Q_{1}z_{k}x_{b}Q_{b}y_{b}z_{t}^{\prime}Q_{2}z_{1}^{\prime}z_{1}^{\prime\prime}Qyy_{a}Q_{a}x_{a}z_{1}$ is a transversal directed Hamilton cycle in $\mathcal{D}$. ∎
Assume that $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is odd. Notice that there exists an edge $z_{1}^{\prime}z_{2}^{\prime}\in E(D_{i_{2}}[A\setminus V(P^{3}),B\setminus V(P^{3})])$ for some $i_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P^{3})$. This edge together with the path $P^{3}$ form two rainbow directed paths that satisfy all conditions in Claim 6.10. It follows that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction. Hence $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even.
@@
Since $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$ and $|A|\leq|B|$, each vertex in $A$ has an in-neighbor in $B$ in each $D_{i}$ with $i\in\mathcal{C}_{1}\cup{\rm col}(M)$. Suppose that there exists an edge $y_{2}y_{1}\in E(D_{i_{3}}[B,A])$ for some $i_{3}\in\mathcal{C}_{1}\cup{\rm col}(M)$, where $y_{1}\in A\setminus(V(P^{3})\cup V_{\rm bad})$ and $y_{2}\in B$. If $i_{3}\in{\rm col}(M)$, then assume ${uv}\in E(M[A])\cap E(D_{i_{3}})$. Applying Lemma 6.3 (i)-(ii), we successively connect $P^{2}$, all rainbow edges in MuvM-uv and all rainbow edges in $M^{\prime}$. This results in a rainbow directed path $\tilde{P}^{3}$ with endpoints $\tilde{u}_{1}\in A\setminus V_{\rm bad}$ and $\tilde{w}_{3}\in B\setminus V_{\rm bad}$. Therefore, $P^{3}$ can be written as $uvv^{\prime}\tilde{u}_{1}\tilde{P}^{3}\tilde{w}_{3}$, where the colors of $vv^{\prime}$} and $v^{\prime}\tilde{u}_{1}$} are in $\mathcal{C}_{2}$. Since $|\mathcal{C}_{2}\setminus{\rm col}(P^{3})|$ is even, one has $|\mathcal{C}_{2}\setminus{\rm col}(\tilde{P}^{3})|$ is even. Hence one may let $P^{3}:=\tilde{P}^{3}$ if $i_{3}\in{\rm col}(M)$ and further assume $i_{3}\not\in{\rm col}(P^{3})$.
@@
If $y_{2}\in X^{\prime}$, then by adjusting the selections of $y_{2}^{1},y_{2}^{2}$ and $j_{1},j_{2}$, one may get a rainbow directed path $P^{3}$ such that $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd. This situation therefore reduces to the case where $|\mathcal{C}_{2}\setminus{\rm col}({P^{3}})|$ is odd.
@@
7 Proof of Theorem 5.3
In this section we present the proof of Theorem 5.3. Since the overall strategy parallels that of Theorem 5.2, the argument itself is relatively short; therefore we do not separate out preliminary claims but instead give a complete, self-contained proof of Theorem 5.3.
@@
Choose constant $\epsilon,\delta$ such that $\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1,$ let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. For each $i\in[m]$, since $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal, one may assume that $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ is a $\epsilon$-characteristic partition of $D_{i}$. By Lemma 5.1, one has $|C_{1}^{k}\triangle C_{i}^{k}|\leq 2\delta n$ for all $k\in[4]$ and all $i\in[m]$. Expand $C_{1}^{1}\cup C_{1}^{2}\cup C_{1}^{3}\cup C_{1}^{4}$ into a partition $C^{1}\cup C^{2}\cup C^{3}\cup C^{4}$ of $V$ such that $|C^{1}|-|C^{3}|\leq 1$ and $|C^{2}|=|C^{4}|$. Hence, for $i\in[2,m]$ and $k\in[4]$, we have $|C^{k}\triangle C_{i}^{k}|\leq 2\delta n+2{\epsilon}n<3\delta n$.